# Solving the Optimal Experiment Design Problem with Mixed-Integer Convex Methods

### v5

@@ 5.0.2 Results
Figure 1: The accumulated number of instances solved to optimality over time for the A-Optimal Problem and D-Optimal Problem with both data sets. The upper plot shows the number of instances solved on the problem with independent data and the bottom the number of solved instances on the correlated data.
(a) A-Fusion Problem
(b) D-Fusion Problem
@@
Figure 2: The accumulated number of instances solved to optimality over time for the A-Fusion Problem and D-Fusion Problem with both data sets. The upper plot shows the number of instances solved on the problem with independent data and the bottom the number of solved instances on the correlated data.
Notice that there is a sharp increase in the beginning, especially for Co-BnB and Boscia. In contrast to Boscia, the curves for Co-BnB flatten more after the first 10 seconds. The notable exception is the A-Fusion Problem with correlated data for which the Co-BnB method works consistently better than boscia, see also Table 1. The better performance of Co-BnB can be explained by the different objectives used. For the A-criterion, Co-BnB uses $\log\left(\operatorname{Tr}(X^{-1})\right)$ instead of $\operatorname{Tr}(X^{-1})$ which both Boscia and SCIP utilize. Computing the Optimal Problems and Fusion Problems using the logarithm formulation with both Boscia and SCIP yields twice as good a performance on the A-Fusion problem with correlated data, see Table 2. Observe that the performance of Boscia on the Optimal Problems decreases when $\log\left(\operatorname{Tr}(X^{-1})\right)$ is used as the objective. It should be noted that whether generalized self-concordance holds for the logarithm of the trace inverse (potentially limited to a subset of its domain) is still an open question.
As previously stated, the Co-BnB framework was developed for the case where the number of parameters $n$, and consequently the number of allowed experiment N, is significantly smaller than the number of experiments $m$. There are no advanced Branch-and-Bound algorithmic strategies proposed, such as improved traverse strategies or branching decisions. A greater value of $N$ naturally increases the size of the tree and the number of nodes to be processed. Boscia has the advantage here since it finds many integer feasible points while solving the relaxations which have the potential to improve the incumbent. A better incumbent, in turn, lets us prune non-improving nodes early on.
Observe that the curves for the two Outer Approximation approaches also flatten out. In addition, their increase at the start is not as sharp as for the two Branch-and-Bound approaches. Keep in mind that while the relaxations of the two Branch-and-Bound approaches keep the simple feasible region intact, the Outer Approximation methods add many additional constraints, i.e. the cuts. This results in larger MIPs to be solved. Furthermore, these cuts are dense leading to further difficulty for the MIP solvers.
@@
Aside from the performance comparison of the solvers, we investigate how the problems themselves compare to each other and if the difficulty of the instances is solver-dependent or if there are clear trends.
@@
It can be observed in Table 1 that most solvers solve fewer instances under the A-criterion. The notable exception is Pajarito on the A-Optimal Problem where it solves more instances to optimality compared to the D-criterion. It should be noted, however, that Pajarito encountered slow progress for many instances under the A-criterion because too many cuts were added or the cuts were too close, i.e. the normal vectors of the hyperplane were too parallel to each other.
@@
Taking a look at some example contour plots shown in Figure 3, we observe that the contour lines for the A-criterion are steeper than those of the D-criterion for both the Optimal Problem and the Fusion Problem, respectively. This points to the condition number increasing.
@@
In terms of the data, one could assume that all problems would be easier to solve with independent data. Noticeably, this is not the case. Rather, it differs for the two problem types. The Fusion Problems are easier with independent data, the Optimal Problems are more often solved with the correlated data. Figures 4(a) to 4(d) depict the progress of the incumbent, lower bound, and dual gap within Boscia for selected instances of each combination of problem and data type.
@@
Interestingly, the independent data leads to proof of optimality, i.e. the optimal solution is found early on and the lower bound has to close the gap, regardless of the problem, see Figures 4(a) and 4(c). The difference in problems has, however, an impact on how fast the lower bound can catch up with the incumbent. Comparing Figures 4(a) and 4(c), the lower bound curve in Figure 4(c) gets closer to the incumbent initially, i.e. in the first 2 seconds, and flattens more compared to the lower bound curve in Figure 4(a). The reason for this difference is likely that in the case of the Optimal Problem there is a larger region around the optimal solution where the corresponding points/designs $\mathbf{x}$ provide roughly the same information. These other candidates have to be checked to ensure the optimality of the incumbent and thus the solving process slows down. If we can prove either strong convexity or sharpness for the objectives, this can potentially be used to speed up the progression of the lower bound. Currently, Boscia can only utilize strong convexity but the adaptation for sharpness is in development.
@@
On the other hand, the correlated data leads to solution processes that are very incumbent-driven, i.e. most improvement on the dual gap stems from the improving incumbent, not from the lower bound, as seen in Figures 4(b) and 4(d). Incumbent-driven solution processes can be identified by the dual gap making sudden jumps and the absence of (a lot of) progress between these jumps. As before, the solution process speed depends on the Problem. In Figure 4(b), the dual gap makes big jumps throughout most of the solving process, in contrast to the dual gap Figure 4(d). For both problems, this indicates that the optimal solution is in the interior. In the case of the Fusion Problem, the lower bound is increasing slower compared to the Optimal Problem in Figure 4(b). This points to the existence of many points yielding a similar objective value. The key ingredients for improvement will be the incorporation of more sophisticated primal heuristics and further investigation on how to improve the lower bound progress.
(a) $m=60$ and $n=15$
(b) $m=100$ and $n=10$
(c) $m=100$ and $n=10$
(d) $m=80$ and $n=20$
Figure 4: Progress of the incumbent and the lower bound on the left and progress of the dual gap for (a) the A-Optimal Problem with independent data, (b) the D-Optimal with correlated data and the A-Fusion Problem with (c) independent and (d) correlated data.
@@
Table 1: Comparing the performance of Boscia, Pajarito, SCIP and Co-BnB on the different problems and the different data sets, i.e. A-Fusion (AF), D-Fusion (DF), A-Optimal (AO) and D-Optimal (DO). One data set contains independent data, the other has correlated data.
The average time is taken using the geometric mean shifted by 1 second. Also, note that this is the average time over all instances in that group, i.e. it includes the time outs.
Boscia Co-BnB Pajarito SCIP+OA
Type Corr. Solved after (s) # inst. % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s)
AO no 0 50 58 % 208.53 42 % 640.68 14 % 1901.7
10 37 43 % 973.13 22 % 2284.39 0 % 3600.05
100 30 30 % 2084.72 3 % 3531.73 0 % 3600.07
1000 24 12 % 3221.62 0 % 3602.71 0 % 3600.07
2000 22 5 % 3548.1 0 % 3602.94 0 % 3600.07
AO yes 0 50 82 % 98.5 50 % 541.22 20 % 1591.74
10 37 76 % 330.67 41 % 1145.77 11 % 2777.96
100 23 61 % 1359.18 9 % 3441.78 0 % 3600.05
1000 13 31 % 2922.48 0 % 3600.16 0 % 3600.02
2000 10 10 % 3596.95 0 % 3600.19 0 % 3600.02
AF no 0 50 80 % 54.82 78 % 82.96 12 % 2006.81 38 % 464.82
10 31 68 % 365.69 65 % 543.73 0 % 3600.03 6 % 3153.28
100 21 52 % 1277.13 48 % 1596.99 0 % 3600.04 0 % 3600.07
1000 13 23 % 3081.0 15 % 3514.57 0 % 3600.05 0 % 3600.1
2000 10 0 % 3600.13 10 % 3573.82 0 % 3600.06 0 % 3600.12
AF yes 0 50 26 % 1359.35 50 % 370.57 20 % 1132.66 14 % 1471.59
10 39 13 % 2684.24 36 % 1421.25 0 % 3600.06 3 % 3379.67
100 35 9 % 3123.03 29 % 2225.95 0 % 3600.07 0 % 3600.03
1000 29 0 % 3600.06 14 % 3393.85 0 % 3600.08 0 % 3600.04
2000 28 0 % 3600.06 11 % 3467.38 0 % 3600.08 0 % 3600.04
DF no 0 50 94 % 3.32 86 % 38.28 14 % 1576.5 50 % 333.25
10 6 50 % 858.17 17 % 3336.0 0 % 3613.96 0 % 3600.21
100 5 40 % 1408.69 0 % 3602.04 0 % 3614.84 0 % 3600.23
DF yes 0 50 60 % 50.68 54 % 185.07 14 % 1761.18 28 % 753.56
10 24 17 % 2534.26 8 % 2944.52 0 % 3605.93 0 % 3600.05
100 23 13 % 3151.98 4 % 3458.68 0 % 3606.19 0 % 3600.05
1000 22 9 % 3416.05 5 % 3452.39 0 % 3606.45 0 % 3600.05
D no 0 50 74 % 81.07 58 % 442.28 24 % 732.57
10 33 61 % 479.16 36 % 2199.09 6 % 1182.22
100 21 43 % 1471.23 10 % 3525.97 5 % 1083.49
1000 6 33 % 2691.04 0 % 3603.88 0 % 2735.7
D yes 0 50 100 % 1.26 68 % 223.34 10 % 755.88
10 7 100 % 24.66 0 % 3600.22 0 % 492.01
Table 2: Comparing the performance of Boscia, Pajarito, SCIP and Co-BnB on A-Optimal and A-Fusion problem with both data sets, i.e. independent and correlated data.
In contrast to Table 1, Boscia and SCIP also have $\log\left(\operatorname{Tr}(X^{-1})\right)$ as objective function.
The average time is taken using the geometric mean shifted by 1 second. Also, note that this is the average time over all instances in that group, i.e. it includes the time outs.
Boscia Co-BnB Pajarito SCIP+OA
Type Corr. Solved after (s) # inst. % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s)
AO no 0 50 56 % 325.09 42 % 640.68 14 % 1901.7
10 40 45 % 952.21 28 % 1789.44 0 % 3600.05
100 33 33 % 2127.44 12 % 2986.04 0 % 3600.06
1000 26 15 % 3202.8 0 % 3602.5 0 % 3600.06
2000 24 8 % 3472.93 0 % 3602.7 0 % 3600.07
AO yes 0 50 64 % 185.28 50 % 541.22 20 % 1591.74
10 38 53 % 613.98 34 % 1669.08 0 % 3600.05
100 29 38 % 1641.98 14 % 3346.09 0 % 3600.05
1000 19 5 % 3533.58 0 % 3600.11 0 % 3600.04
AF no 0 50 84 % 31.27 78 % 82.96 12 % 2006.81 38 % 477.72
10 25 68 % 270.78 56 % 1074.55 0 % 3600.03 0 % 3600.06
100 14 43 % 1457.98 21 % 2474.08 0 % 3600.05 0 % 3600.08
1000 9 11 % 3553.67 11 % 3570.97 0 % 3600.06 0 % 3600.12
AF yes 0 50 52 % 207.34 50 % 370.57 20 % 1132.66 26 % 934.9
10 33 27 % 1780.79 24 % 2097.49 0 % 3600.07 3 % 3278.79
100 29 17 % 3075.17 21 % 2656.55 0 % 3600.07 0 % 3600.04
1000 26 8 % 3461.56 12 % 3457.38 0 % 3600.07 0 % 3600.05
2000 25 4 % 3545.06 8 % 3513.62 0 % 3600.07 0 % 3600.05
@@ 6 Conclusion
We proposed a new approach for the Optimal Experiment Design Problem based on the Boscia framework, and show its performance compared to other MINLP approaches. In addition, it also performs better compared to the approach specifically developed for the OEDP, in particular for large-scale instances and a larger number of parameters. This superiority can be explained by the fact that Boscia keeps the structure of the problem intact and that it utilizes a combinatorial solver to find integer feasible points at each node.
{credits}
6.0.1 Acknowledgements
@@
Appendix 0.A Convergence proofs on continuous relaxations
0.A.1 Lipschitz smoothness of the Fusion Problems
Definition 5
A function $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is called $L$-smooth if its gradient $\nabla f$ is Lipschitz continuous. That is, there exists a finite $L\in{\mathbb{R}}$ with
$\left\|\nabla f(\mathbf{x})-\nabla f(\mathbf{y})\right\|_{2}\leq L\left\|\mathbf{x}-\mathbf{y}\right\|_{2}.$
Since the functions of interest are in $\mathcal{C}^{3}$, we can use an alternative condition on $L$-smoothness: $LI\succcurlyeq\nabla_{xx}f(\mathbf{x})$, i.e. the Lipschitz constant $L$ is at least as large as the maximum eigenvalue of the Hessian.
We will see that we have $L$-smoothness on the whole convex feasible region $\mathcal{P}$ only for the Fusion Problems. Due to the positive definiteness of the matrix $C$, the information matrix in the Fusion case $X_{C}(\mathbf{x})$ is always bounded away from the boundary of the PSD cone. As for the Optimal Problems, we will prove self-concordance in Section 0.A.2 to consequently show convergence for Frank-Wolfe.
Before we investigate $L$-smoothness, we first prove a relationship between a matrix $X$ and its power matrix $X^{r}$ where $\0r\in{\mathbb{R}}\backslash\{0\}$.
Lemma 1
Let $X\in{\mathbb{R}}^{n\times n}$ by an invertible matrix and $\0r\in{\mathbb{R}}\backslash\{0\}$. Then, the eigenvalues of the power matrix are:
$\lambda_{i}(X^{r})=\lambda_{i}(X)^{r}\quad\forall i=[n]$
and the eigenvectors of the power matrix are the same as those of $X$.
Proof
For any square matrix $A$ and a positive integer $m$, we have that $A^{m}\mathbf{v}=\lambda_{\mathbf{v}}(A)^{m}\mathbf{v}$. Also, the logarithm of a matrix $A$ has the same eigenvectors as the original and its eigenvalues are $\log(\lambda_{i}(A))$.
$=\lambda_{\mathbf{v}}(X)^{r}\mathbf{v}$
This concludes the proof. ∎
Additionally, we need the following statement about symmetry.
Lemma 2
@@ 0.A.1 Lipschitz smoothness of the Fusion Problems
Corollary 1
For $\0r\in{\mathbb{R}}\backslash\{0\}$, if $X\in\mathbb{S}_{++}^{n}$, then $X^{r}\in\mathbb{S}_{++}^{n}$.
In order to compute the derivatives of the functions in question, we require the following lemma.
Lemma 3
Let $r\in{\mathbb{R}}$ and let the matrix $X$ be diagonalizable. Then, we can define the function $f(X)=\operatorname{Tr}(X^{r})$. The gradient of $f$ is then
@@
Thus, using the sum rule, $(\operatorname{Tr}(X^{n}))^{\prime}=nX^{n-1}$ for any positive integer $n$ and $\nabla\log(X)=\nabla XX^{-1}$, we have
$\nabla\operatorname{Tr}(X^{r})$ $=\sum\limits_{k=1}^{\infty}\frac{r^{k}}{k!}k{\log(X)}^{k-1}IX^{-1}$
$=r\left(\sum\limits_{k=0}^{\infty}\frac{r^{k}}{k!}\log(X)^{k}\right)X^{-1}$
$=rX^{r}X^{-1}=rX^{r-1}$

### v6

@@ Results.
Figure 4: The number of instances solved to optimality over time for the A-Optimal Problem and D-Optimal Problem with both data sets. The upper plots are with the inpdendent data set, the lower ones with correlated data. Note that SCIP+OA is not applicable on the Optimal Problems.
(a) A-Fusion Problem with independent data
(b) D-Fusion Problem with independent data
(c) A-Fusion Problem with correlated data
(d) D-Fusion Problem with correlated data
@@
Figure 5: The number of instances solved to optimality over time for the A-Fusion Problem and D-Fusion Problem with both data sets. The upper plots are with the independent data set, the lower ones with correlated data.
For the (GTI-Opt) Problem, the results are given in Table 3. Note that we used Boscia.jl on both the (GTI-Opt) and the (logGTI-Opt), the latter titled as "Boscia Log" in the table. Boscia.jl also shows a better performance here than Co-BnB and SCIP + OA. Interestingly, the (logGTI-Opt) objective seems to lead to a better conditioning of the problem. This further motivates an investigation into generalized self-concordance of those objectives.
In Table 4, we see the relative and absolute gap of Boscia.jl and Co-BnB for instances with sizes from 300 and 500 variables. Boscia.jl showcases better gap values especially for the problems under the D-criterion. Note that almost none of the instances were solved to optimality.
We performed a hyperparameter search for Boscia.jl the results of which can be found in Tables 8, 9 and 10.
@@
Aside from the performance comparison of the solvers, we investigate how the problems themselves compare to each other and if the difficulty of the instances is solver-dependent.
@@
It can be observed in Table 2 that most solvers solve fewer instances under the A-criterion. The notable exception is SOCP on the A-Fusion Problem where it solves more instances to optimality compared to the D-criterion.
@@
Taking a look at some example contour plots shown in Figure 6, we observe that the contour lines for the A-criterion are steeper than those of the D-criterion for both the Optimal Problem and the Fusion Problem, respectively. This points to the conditioning increasing.
@@
In terms of the data, one could assume that all problems would be easier to solve with independent data. Noticeably, this is not the case, rather, it depends on the problem type. The Fusion Problems are easier to solve with independent data, the Optimal Problems are more often solved with the correlated data. Figures 7(a) to 7(d) depict the progress of the incumbent, lower bound, and dual gap within Boscia.jl for selected instances of each combination of problem and data type.
@@
Interestingly, the independent data leads to proof of optimality, i.e. the optimal solution is found early on and the lower bound has to close the gap, regardless of the problem, see Figures 7(a) and 7(c). The difference in problems has, however, an impact on how fast the lower bound can catch up with the incumbent. In the case of the Optimal Problem there is likely a larger region around the optimal solution where the corresponding points/designs $\mathbf{x}$ provide roughly the same information. These other candidates have to be checked to ensure the optimality of the incumbent and thus the solving process slows down. Note that Boscia.jl can utilize strong convexity and sharpness of the objective to improve the lower bound. However, the improvement via sharpness is currently only applicable if the optimum is unique which is not the case.
@@
On the other hand, the correlated data leads to solution processes that are very incumbent-driven, i.e. most improvement on the dual gap stems from the improving incumbent, not from the lower bound, as seen in Figures 7(b) and 7(d). Incumbent-driven solution processes can be identified by the dual gap making sudden jumps and the absence of (a lot of) progress between these jumps. As before, the solution process speed depends on the Problem. In Figure 7(b), the dual gap makes big jumps throughout most of the solving process, in contrast to the dual gap Figure 7(d). This indicates that the optimal solution of the Fusion Problem is strictly in the interior and thus, will not be found early as a vertex of a relaxation. The key ingredient for improvement in this case will be the incorporation of more sophisticated primal heuristics in Boscia.jl.
Overall, Boscia.jl shows a very promising performance and further research and development will be performed to improve the performance even further.
(a) D-Optimal Problem with independent data and $n=10$
(b) D-Optimal Problem with correlated data and $n=10$
(c) A-Fusion Problem with independent data and $n=10$
(d) A-Fusion Problem with correlated data and $n=25$
Figure 7: Progress of the incumbent and the lower bound on the left and progress of the absolute gap on the right. All instances have 100 variables.
@@
Table 2: Comparing the performance of Boscia.jl, Pajarito.jl, SCIP, Co-BnB and SOCP on the different problems and the different data sets, i.e. A-Fusion (AF), D-Fusion (DF), A-Optimal (AO) and D-Optimal (DO). One data set contains independent data, the other has correlated data. Shown are the percentage of solved instances and the geometric mean of the solve time (shifted by 1 second) over all instances. Note that there are 50 instances for each problem and data set.
For more extensive results like number of cuts/nodes, see Table 5 in Section D.
Boscia Co-BnB Direct Conic SOCP SCIP OA
Type Corr. % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s)
AO no 52 % 374.74 48 % 401.67 8 % 2382.52 20 % 1578.03 - -
AO yes 68 % 227.69 60 % 301.8 20 % 1515.23 20 % 1861.35 - -
AF no 80 % 62.83 80 % 50.92 4 % 2846.5 26 % 1686.2 38 % 463.65
AF yes 26 % 1368.83 52 % 311.36 24 % 1148.87 20 % 2131.81 14 % 1480.73
DO no 68 % 136.92 58 % 216.31 20 % 1241.37 20 % 2169.4 - -
DO yes 100 % 2.56 70 % 101.89 10 % 1398.01 10 % 2526.74 - -
DF no 94 % 3.49 86 % 22.79 12 % 1906.61 18 % 2213.03 50 % 324.11
DF yes 60 % 54.01 56 % 153.26 6 % 2284.06 6 % 2934.61 28 % 752.15
Table 3: Comparing the performance of Boscia.jl,SCIPand Co-BnB on the General-Trace-Inverse (GTI) Optimal Problem and GTI-Fusion (GTIF) Problem for different values of $p$ and under independent and correlated data. Note the the objective function of Boscia has the form $\phi_{p}(X(\mathbf{x}))=\operatorname{Tr}\left(X(\mathbf{x})^{-p}\right)$. The column titled "Boscia Log" shows the runs using $\log\phi_{p}$ as the objective. Shown are the percentage of solved instances and the geometric mean of the solve time (shifted by 1 second) over all instances. Note that there are 30 instances for each problem and data set.
For extensive results, see Tables 6 and 7 in Section D.
Boscia Boscia Log Co-BnB SCIP+OA
p Corr. % solved Time (s) % solved Time (s) % solved Time (s) % solved Time (s)
Optimal Problem
0.25 no 67 % 162.41 87 % 27.13 50 % 359.92 - -
0.25 yes 67 % 110.25 77 % 85.83 60 % 264.22 - -
0.5 no 53 % 246.61 70 % 85.87 50 % 430.08 - -
0.5 yes 60 % 129.86 57 % 245.36 57 % 313.7 - -
0.75 no 37 % 799.73 43 % 505.86 50 % 471.95 - -
0.75 yes 47 % 541.67 43 % 557.15 50 % 357.77 - -
1.5 no 43 % 659.25 40 % 576.51 50 % 619.95 - -
1.5 yes 47 % 418.43 53 % 306.7 47 % 506.3 - -
2.0 no 43 % 661.29 50 % 562.4 50 % 630.18 - -
2.0 yes 43 % 644.53 53 % 297.98 47 % 611.8 - -
Fusion Problem
0.25 no 90 % 6.57 100 % 1.71 50 % 725.52 47 % 376.08
0.25 yes 53 % 138.78 57 % 184.83 27 % 1449.73 27 % 927.92
0.5 no 87 % 9.1 90 % 3.63 47 % 787.45 43 % 349.14
0.5 yes 57 % 175.85 53 % 271.12 30 % 1547.49 20 % 850.41
0.75 no 70 % 85.39 87 % 42.47 57 % 353.3 43 % 415.34
0.75 yes 37 % 741.57 43 % 434.78 27 % 1639.63 23 % 1053.8
1.5 no 77 % 91.83 67 % 129.57 47 % 780.56 40 % 406.94
1.5 yes 37 % 715.74 50 % 214.97 30 % 1415.97 23 % 885.83
2.0 no 73 % 96.28 70 % 117.1 57 % 797.56 40 % 402.28
2.0 yes 33 % 905.59 53 % 202.0 33 % 1305.28 20 % 923.45
Table 4: Comparing Boscia and Co-BnB on large instances. The dimensions are $m=300,400,500$. We showcase only the absolute and relative gap. Note that the majority of the instances were solved to optimality. Boscia solved 3 instances of the D-Fusion Problem with independent data. The relative gap and the absolute gap are computed for the instances on which at least one method did not terminate within the time limit. That means it excludes the instances on which all methods terminated. The "log" prefix indicates that Boscia was used with the $\log\operatorname{Tr}$ objective.
Boscia Co-BnB
Problem Corr. # inst. relative gap absolute gap relative gap absolute gap
A no 12 188.2483 3.8117 2.4036 5.2537
A yes 12 Inf 0.0024 Inf 5.8384
AF no 12 0.2243 0.0019 0.9264 1.0258
AF yes 12 Inf 0.0049 Inf 5.6742
log A no 12 2.4047 5.2608 2.4036 5.2537
log A yes 12 Inf 5.6258 Inf 5.8384
log AF no 12 1.0491 2.6105 0.9264 1.0258
log AF yes 12 0.0955 6.9323 Inf 5.6742
DF no 12 0.0169 0.01 Inf 38.0805
DF yes 12 0.9819 0.4054 Inf 343.1067
D no 12 0.2295 0.1577 Inf 69.0189
D yes 12 0.4642 0.091 673.4528 73.5618
@@ 6 Conclusion
We proposed a new approach for the Optimal Experiment Design Problem based on the Boscia.jl framework and proved convergence of the method on the problems. Our approach exhibits a superior performance compared to other MINLP approaches. In addition, it also outperforms the approach specifically developed for the OEDP, in particular for large-scale instances and a larger number of parameters. This superiority can be explained by the fact that Boscia.jl keeps the structure of the problem intact and that it utilizes a combinatorial solver to find integer feasible points at each node.
Acknowledgments
@@
A Additional proofs and calculations
Note that the following lemmas and the proofs are not new but they are stated here for completeness.
A.1 General power of a matrix
To start, we present some smaller statements about the power of a matrix. These are used in the consequent proofs.
Lemma A.17.
@@ A.1 General power of a matrix
Lemma A.18.
Let $X\in{\mathbb{R}}^{n\times n}$ by an invertible matrix and $r\in{\mathbb{R}}$. Then, the eigenvalues of the power matrix are
$\lambda_{i}(X^{r})=\lambda_{i}(X)^{r}\quad\forall i=[n]$
and the eigenvectors of the power matrix are the same as those of $X$.
Proof.
For any square matrix $A$ and a natural number $m\in{\mathbb{N}}$, we have that $A^{m}\mathbf{v}=\lambda_{\mathbf{v}}(A)^{m}\mathbf{v}$. Also, the logarithm of a matrix $A$ has the same eigenvectors as the original and its eigenvalues are $\log(\lambda_{i}(A))$.
$=\lambda_{\mathbf{v}}(X)^{r}\mathbf{v}$
This concludes the proof. ∎
Combining Lemma A.17 and Lemma A.18, we get the following corollary.
Corollary A.19.
For $r\in{\mathbb{R}}$, if $X\in\mathbb{S}_{++}^{n}$, then $X^{r}\in\mathbb{S}_{++}^{n}$.
Lemma A.20.
Let $r\in{\mathbb{R}}$ and let the matrix $X$ be diagonalizable. Let $f(X)=\operatorname{Tr}(X^{r})$. The gradient of $f$ is then
@@
Thus, using the sum rule, $(\operatorname{Tr}(X^{n}))^{\prime}=nX^{n-1}$ for any positive integer $n$ and $\nabla\operatorname{Tr}(\log(X))=X^{-1}$, we have
$\nabla\operatorname{Tr}(X^{r})$ $=\sum\limits_{k=1}^{\infty}\frac{r^{k}}{k!}k{\log(X)}^{k-1}X^{-1}$
$=rX^{-1}\sum\limits_{k=0}^{\infty}\frac{r^{k}}{k!}\log(X)^{k}$
$=rX^{-1}X^{r}=rX^{r-1}$